a mass of 7.50 kg is being pushed down a vertical rough wall by a force F=250 N at angel 30 degrees from horizontal. the kinetic coefficient of friction between wall and mass is 0.430.

A. Find frictional force.
B. find acceleration of mass
C, if mass move along wall 3.2 Meters - find work done only by F force
Answer #1

The normal reaction force from the surface of the wall onto the block is

N = F \cos 30 = 250 \cos 30 = 216.5N

The frictional force is

f = \mu_k N = (0.43)(216.5) = 93.1N

b)

the accleration of the mass

mg + F \sin 30 - f = m a

(7.5)(9.8) + (250) \sin 30 - 93.1 = (7.5) a

The acceleration is

a = 14 m/s^2

c) The work done only by the force is

W = F \sin 30 . d

W = (250) \sin 30 . (3.2) = 400 J